# Geometric insights into robust portfolio construction

### v4

@@ 6 Conclusions
Mean-variance portfolio decisions that combine prediction and optimisation have been shown to have poor empirical performance for a variety of shrinkage predictors under different distributional assumptions that combined mean-variance optimisation with reasonable departures from Gaussianity [40]. Here we directly consider both constrained and unconstrained mean-variance portfolio construction and performance in a general setting that does not have strong distributional assumptions, and we confirm these theoretical results using a simple Gaussian simulation. This is made possible by casting mean-variance portfolio theory into a geometric setting.
To present the geometric perspective we consider the upper bound of the $\boldsymbol{\alpha}$-weight angle as introduced by Golts and Jones [19]. First, we show how this result arises as a result of the Kantorovich inequality in Theorem 1. Despite being framed in the mean-variance setting the $\boldsymbol{\alpha}$-angle approach to covariance shrinkage makes no distributional assumptions about the data generating processes. Golts and Jones [19] only considered unconstrained optimisation problems when demonstrating the relationship between the $\boldsymbol{\alpha}$-weight angle and the condition of the covariance matrix; as in optimisations I, II, III and IV in Table 1. Here we show this result can be weakened and generalised when gearing constraints are added; as in the optimisations V, VI, VII and VIII in Table 2 using the extension of Bauer and Householder given by Theorem 2.
@@
This is demonstrated for the case of normal returns in Section 5. Where we show that in the case of niave unconstrained portfolio optimisation that the equally-weighted portfolio’s performance better than the mean-variance optimal portfolios. However, when applying the method first proposed by [19] this is no longer the case, the mean-variance portfolio’s would on average out-perform the equally weighted portfolio’s. Similarly in the presence of linear constraints, when the shrinkage hyper-parameters are appropriate selected one can ensure that the constrained mean-variance portfolio’s, on average, out-perform the equally weighted portfolio’s. This is consistent with the distribution free general result used to motivate the method of [19].
@@ Appendix B Bauer and Householder Inequality
If $\widetilde{M}$ has characteristic roots $\tau$ and $\tau\kappa(\widetilde{M})$ then $\det\widetilde{M}=\tau^{2}\kappa(\widetilde{M})$ and $(\widetilde{m}_{11}\widetilde{m}_{22})^{\frac{1}{2}}\leq\frac{1}{2}(\widetilde{m}_{11}+\widetilde{m}_{22})=\frac{\tau}{2}(1+\kappa(\widetilde{M}))$ so that 58 is bounded above
@@
Using that $\mathbf{x}$ is orthogonal to $Y$ where the cosine of the angle between $\mathbf{r}$ and the subspace of $Y$ is $\cos(Y,\mathbf{r})={\|Y(Y^{{}_{H}}Y)^{-1}Y\mathbf{r}\|}/{\|\mathbf{r}\|}$ we have that $\cos(\mathbf{x},\mathbf{r})\leq[1-\cos^{2}(Y,\mathbf{r})]^{\frac{1}{2}}$. Then we have that $\|Y(Y^{{}_{H}}Y)^{-1}Y^{{}_{H}}\mathbf{r}\|\geq\|\mathbf{r}\|\sin\psi$ which implies that $\|H\mathbf{s}\|\leq\|\mathbf{s}\|\cos\psi$. From this an optimal bound for 61, and thus 62, can be obtained when the space of $Y$ contains $\mathbf{r}$:

### v5

@@ 6 Conclusions
Mean-variance portfolio decisions that combine prediction and optimisation have been shown to have poor empirical performance for a variety of shrinkage predictors under different distributional assumptions that combined mean-variance optimisation with reasonable departures from Gaussianity [44]. Here we directly consider both constrained and unconstrained mean-variance portfolio construction and performance in a general setting that does not have strong distributional assumptions and we then confirm these theoretical results using a simple Gaussian simulation. This is made possible by casting mean-variance portfolio theory into a geometric setting.
To present the geometric perspective we consider the upper bound of the $\boldsymbol{\alpha}$-weight angle as introduced by Golts and Jones [19]. First, we show how this result arises as a result of the Kantorovich inequality in Theorem 1. Despite being framed in the mean-variance setting the $\boldsymbol{\alpha}$-angle approach to covariance shrinkage makes no distributional assumptions about the data generating processes. Golts and Jones [19] only considered unconstrained optimisation problems when demonstrating the relationship between the $\boldsymbol{\alpha}$-weight angle and the condition of the covariance matrix; as in optimisations I, II, III and IV in Table 1. Then, we show that this result can be weakened and generalised when gearing constraints are added; as in the optimisations V, VI, VII and VIII in Table 2 using the extension of Bauer and Householder given by Theorem 2.
@@
This is demonstrated for the case of normal returns in Section 5.2. Where we show that in the case of niave unconstrained portfolio optimisation that the equally-weighted portfolio’s performance better than the mean-variance optimal portfolios. However, when applying the method first proposed by [19] this is no longer the case, the mean-variance portfolio’s would on average out-perform the equally weighted portfolio’s. Similarly in the presence of linear constraints, when the shrinkage hyper-parameters are appropriate selected one can ensure that the constrained mean-variance portfolio’s, on average, out-perform the equally weighted portfolio’s. This is consistent with the distribution free general result used to motivate the method of [19].
This is in turn confirmed using historic simulation in Section 5.3. Here we find that under unconstrained solutions and constrained solutions V, VII and VIII that the method of Golts and Jones [19] delivers portfolios that are better aligned to the returns of the assets, and generating out-of-sample performance with superior Sharpe ratios than the other methods considered. However, we note that under constrained solution VI, this is not the case because that particular solution naturally tends to solve for weights that generate that minimum returns exactly, and as such moving the portfolio weights closer to the expected return vector is not beneficial despite improve the conditioning of the covariance matrix.
@@ Appendix B Bauer and Householder Inequality
If $\widetilde{M}$ has characteristic roots $\tau$ and $\tau\kappa(\widetilde{M})$ then $\det\widetilde{M}=\tau^{2}\kappa(\widetilde{M})$ and $(\widetilde{m}_{11}\widetilde{m}_{22})^{\frac{1}{2}}\leq\frac{1}{2}(\widetilde{m}_{11}+\widetilde{m}_{22})=\frac{\tau}{2}(1+\kappa(\widetilde{M}))$ so that 59 is bounded above
@@
Using that $\mathbf{x}$ is orthogonal to $Y$ where the cosine of the angle between $\mathbf{r}$ and the subspace of $Y$ is $\cos(Y,\mathbf{r})={\|Y(Y^{{}_{H}}Y)^{-1}Y\mathbf{r}\|}/{\|\mathbf{r}\|}$ we have that $\cos(\mathbf{x},\mathbf{r})\leq[1-\cos^{2}(Y,\mathbf{r})]^{\frac{1}{2}}$. Then we have that $\|Y(Y^{{}_{H}}Y)^{-1}Y^{{}_{H}}\mathbf{r}\|\geq\|\mathbf{r}\|\sin\psi$ which implies that $\|H\mathbf{s}\|\leq\|\mathbf{s}\|\cos\psi$. From this an optimal bound for 62, and thus 63, can be obtained when the space of $Y$ contains $\mathbf{r}$: